OpenStudy (anonymous):

does anyone know how this would be solved? e^xe^(x+1)=1

OpenStudy (agl202):

e^(1.609) = 4.998 (positive because it was -x and the answer was negative) regardless, to solve an equation like this you need to know inverse operations: the inverse operation for e is ln (natural log), because ln(e^x) = x and e^ln(x) = x therefore, you would solve the equation by taking the ln of both sides ln(e^-x) = ln(4) therefore, -x = ln(4) = 1.386, so x = -1.386 that would be the answer unless i am misunderstanding the syntax of the math problem. Hope this helped! :)

OpenStudy (alekos):

start by using the exponents rule a^x * a^x = a^2x

OpenStudy (anonymous):

\[e ^{xe ^{x+1}}=1\]

OpenStudy (anonymous):

its e^(x) e^(x+1)=1 so not with the exponent having another exponent added to it sorry i dont know how to write it correctly :P

OpenStudy (alekos):

yeah I worked it out. just start with what I wrote before

OpenStudy (anonymous):

does that mean that its like this e^x * e^(x+1) = e^2x+1 e^2x+1 =1

OpenStudy (alekos):

yes, that's it. now for the rest we can see that e^0 =1 so e^(2x+1) = e^0 what do you think we do next?

OpenStudy (anonymous):

2X+1=0 2X=-1 X=-1/2

OpenStudy (alekos):

very good! you're getting the hang of this

OpenStudy (anonymous):

thanks :D

OpenStudy (alekos):

no problem

OpenStudy (alekos):

with all of these if you want to check the answer just substitute the final x value into the original equation

OpenStudy (anonymous):

ok cool thank you!!!